OpenStudy (anonymous):

How to add 7 1/6 + 4 1/6?

OpenStudy (solomonzelman):

you can add 7+4 and 1/6 + 1/6 and then reduce the fraction and put the number and fraction to gether to form a mixed number which will be your final answer.

OpenStudy (♪chibiterasu):

\[7 \frac{ 1 }{ 6 } + 4 \frac { 1 }{ 6 } = 11 \frac { 2 }{ 6 }\] Then simplify, the highest common multiple for the fractions is 2 so: \[11 \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

thank you so much guys!